OpenStudy (anonymous):

Multiply and then simplify the radical. 5 square root of 10 times 3 square root of 8

OpenStudy (anonymous):

@PinkSapphire

OpenStudy (pinksapphire):

Can you try it out this time? Then I'll tell you if you did it correctly or not.

OpenStudy (anonymous):

15 square root of 18

OpenStudy (pinksapphire):

I don't think you can do that because when you're multiplying radicals, treat them as if they are variables. For example:\[5x \times 3x=15x\]But for this problem, you have to make the radicals (square roots) the same first before yu multiply the numbers outside the radicals.

OpenStudy (pinksapphire):

*you

OpenStudy (pinksapphire):

Actually, your answer was right, my mistake.

OpenStudy (pinksapphire):

Wait, it's actually...\[15\sqrt{80}\]

OpenStudy (anonymous):

You multiplied it, but you still have to simplify \(15 \sqrt{80}\).

OpenStudy (pinksapphire):

\[5\sqrt{10}\times3\sqrt{8}\] \[5\times3 \sqrt{10\times8}\] \[15\sqrt{80}\]But to simplify it... \[15\sqrt{10\times8}\] \[15\sqrt{5\times2\times4\times2}\] \[15\sqrt{5\times4\times4}\] \[15\sqrt{5\times16}\] \[15\times4\sqrt{5}\] \[60\sqrt{5}\] I'm not really sure if I did it correctly...